OpenStudy (anonymous):

NEED HELP PLEASE:What are the first five terms of the pattern with the formula 2n+2

OpenStudy (sidsiddhartha):

put n=1,2,3,4,5 2(1) + 2 = 4 2(2) + 2 = 6 2(3) + 2 = 8 2(4) + 2 = 10 2(5) + 2 = 12

OpenStudy (anonymous):

wow thanks your smart :)